OpenStudy (anonymous):

Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (anonymous):

Let A = monthly allowance, M = 12A (allowance total for one year) She starts with M, spends 1/2M on clothes and is left with: M - 1/2M = 1/2M = M/2 Next, she spends 1/3 of her remaining money (M/2) on games, leaving her with: M/2 - (1/3)(M/2) = M/2 - M/6 = 2M/6 = M/3 Next she spends 1/4 of her remaining money (M/3) on toys, leaving her with: M/3 - (1/4)(M/3) = M/3 - M/12 = 4M/12 - M/12 = 3M/12 = M/4 This is her remaining money, M/4, which is 7777. So, M/4 = 7777 M = 7777(4) M = 31,108 But we want to know A: M = 12A A = M/12 A = 31,108/12 A = 2592.33, her monthly allowance Let's check: (2592.33)(12) = 31108 Starting money Spend 1/2 on clothes: 31108 - (1/2)(31108) = 15554 Spend 1/3 of this on games: 15554 - (1/3)(15554) = 15554 - 5184.67 = 10369.33 Spend 1/4 of this on toys: 10369.33 - (1/4)(10369.33) = 10369.33 - 2492.33 = 7777

OpenStudy (anonymous):

$2222

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

nope, $2222 doesn't work :-( the right answer is 2592.33